On the 3rd quarter examination in Statistics, the population mean was 70 and the population standard deviation was 9. Determine the standard score of a student who got a score of 88 assuming that the scores are normally distributed.


1
Expert's answer
2022-02-21T07:46:39-0500

If "X\sim N(\mu, \sigma^2)," then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)."

Given "\mu=70, \sigma=9."


"Z=\dfrac{88-70}{9}=2"